OpenStudy (rowanwhitethorn):

The area of a regular hexagon is 40 cm^2. What is the area of a regular hexagon with sides four times as long?

jimthompson5910 (jim_thompson5910):

If the sides are 4 times as long, then the area is 16 times bigger 4^2 = 4*4 = 16

jimthompson5910 (jim_thompson5910):

so you need to multiply the original area (40) by 16 to get your answer

OpenStudy (rowanwhitethorn):

thank you so much!!

jimthompson5910 (jim_thompson5910):

no problem